Q- Find the angle whose complementry angle is 18 more.
A:- is 54 but how??? plz explain.


Answer by Theo(13342)

if an angle is x degrees, then its complementary angle is 90-x because the sum of the angle and its complementary angle must equal 90 degrees.
-----
your problem asks to find the angle who complementary angle is 18 degrees more than the angle.
-----
let your angle = x
let your complementary angle = x+18
-----
since an angle and its complementary angle must equal 90 degrees, your equation is:
x + (x + 18) = 90
remove parentheses to get x + x + 18 which is the same as 2x + 18 = 90
subtract 18 from both sides of this equation and you get 2x = 72
divide both sides of this equation by 2 and you get x = 36
-----
your angle is 36 degrees
your complementary angle is 36 + 18 = 54 degrees
54 + 36 = 90
-----